# DHP Benchmark: Are LLMs Good NLG Evaluators?

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly serving as evaluators in Natural Language Generation (NLG) tasks. However, the capabilities of LLMs in scoring NLG quality remain inadequately explored. Current studies depend on human assessments and simple metrics that fail to capture the discernment of LLMs across diverse NLG tasks. To address this gap, we propose the Discernment of Hierarchical Perturbation (DHP) benchmarking framework, which provides quantitative discernment scores for LLMs utilizing hierarchically perturbed text data and statistical tests to measure the NLG evaluation capabilities of LLMs systematically. We have re-established six evaluation datasets for this benchmark, covering four NLG tasks: Summarization, Story Completion, Question Answering, and Translation. Our comprehensive benchmarking of five major LLM series provides critical insight into their strengths and limitations as NLG evaluators.
@@
Yicheng Wang1*, Jiayi Yuan2*, Yu-Neng Chuang2, Zhuoer Wang1, Yingchi Liu3, Mark Cusick3, Param Kulkarni3, Zhengping Ji 3, Yasser Ibrahim3, Xia Hu2, 1Texas A&M University, 2Rice University , 3Axon Enterprise, Inc., * Equal Contribution
@@ 1 Introduction
Large Language Models (LLMs) play a crucial role in the field of Natural Language Generation (NLG), advanced wide real-world applications including education Latif et al. (2023), healthcare Yuan et al. (2023), business Teubner et al. (2023), etc. The strong capabilities of LLMs allow them not only to serve as text generators but also increasingly as powerful evaluators of text quality Chiang and Lee (2023); Liu et al. (2023a); Li et al. (2024). Their role as evaluators is crucial for advancements in various applications, such as summarization, story completion, question answering, and translation Li et al. (2024); Wang et al. (2023a). LLMs are expected to serve as NLG evaluators, providing reasonable quality scores based on different quality metrics with specially designed evaluation prompts.
@@
(1) Lack of Clear and Unbiased Measurement: There is no clear measurement for the capability of LLM evaluators. Existing methods rely on aligning with human scores Chiang and Lee (2023); Liu et al. (2023a) but these scores themselves are subject to biased response styles Schoch et al. (2020).
(2) Multiple Evaluation Metrics: Evaluating NLG quality requires considering multiple metrics. For example, in summarization tasks, metrics such as coherence, consistency, and fluency are essential considerations Hu et al. (2024); Fabbri et al. (2021). However, LLMs might struggle with correlations between these metrics, potentially leading to misinterpretation and incorrectly scoring, which makes it difficult to assess their effectiveness as evaluators.
To address these challenges, we introduce a novel DHP benchmarking frameworkDiscernment of Hierarchical Perturbation – for quantitatively measuring the evaluation capabilities of LLMs without the need for human annotations. We propose the concept of discernment scores, systematically derived from hierarchically perturbed text data and statistical tests. We perturb the reference data using various hierarchical methods, then compare the differences in LLM evaluation scores using the Wilcoxon Signed-Rank Test Wilcoxon (1945). In order to get overall capability results, we introduce the approach of harmonic mean $p$-values and expert weights to combine the results of multiple metrics. The final $p$-value derived can be transformed to Discernment Scores to measure the NLG evaluation capabilities of LLMs. This method allows for a more rigorous and comprehensive evaluation of LLM performance, independent of the response styles of the tested models.
@@
Our study re-establishes six evaluation datasets that encompass four key NLG tasks: Summarization, Story Completion, Question Answering, and Translation. Each dataset is perturbed and leveraged to challenge the LLMs’ evaluative capabilities in distinct ways, providing a robust basis for benchmarking. The datasets include a range of text perturbations, from minor character problems to significant sentence alterations, allowing us to test the potential discernment limits of the LLMs.
@@
Our comprehensive benchmarking, using newly defined quantitative discernment scores, analyzes five major LLM series. This approach uncovers critical insights into their capabilities as NLG evaluators and provides a detailed understanding of their performance. This benchmark reveals important trends and patterns in the capabilities of LLMs, highlighting areas where they excel and where they may fall short.
@@
The DHP benchmark aims to fill the existing gaps by providing a quantitative framework for assessing LLMs’ evaluation capabilities and emphasizing the necessity of considering multiple metrics for accurate and reliable evaluations. We summarize our contributions as follows.
  1. (1)
@@
    Design the DHP benchmarking framework: quantitative discernment scores for LLMs as NLG evaluators based on hierarchical perturbation, eliminating the need for human annotations.
  2. (2)
@@
    Re-establish six evaluation datasets for four NLG tasks to evaluate the discernment of LLMs.
  3. (3)
@@
    Benchmark five series of LLMs to analyze their capabilities of NLG evaluation.
2 Related Work
Recent advancements highlight the significant potential of utilizing LLMs as evaluators for a variety of natural language processing (NLP) tasks. Extensive empirical evidence supports this viewpoint, as demonstrated by studies Liu et al. (2023a); Chiang and Lee (2023); Hu et al. (2024); Desmond et al. (2024); Wang et al. (2023a), which assert that the evaluation behaviors of pretrained LLM-based evaluators are well-aligned with those of human preference Liu et al. (2023b) and eliminate the prompt bias Liusie et al. (2024). Another line of work Huang et al. (2024); Zhu et al. (2023); Wang et al. (2023b) finetunes LLMs with specific downstream task datasets to increase the capability as the judges. Despite the great assessment performance of a single LLM, advanced studies involve multi LLM agents Chan et al. (2023); Zhang et al. (2023); Li et al. (2023) or human experts Gao et al. (2024); Li et al. (2024) to further increase the judging capability. However, humans often prefer to remain neutral Leroi et al. (2020), which means a perfect alignment between LLM-based evaluators and human evaluators may still lead to bias and inaccurate judgments.
@@ 2 Related Work
While using LLM-based evaluators offers a scalable and effective method for approximating human preferences, the significant influence of individual human biases raises concerns about their discernment in assessing LLM-based evaluations. In our work, we focus on developing comprehensive approaches to observe and address this phenomenon.
3 Biased Response Styles
@@
We illustrate this with an example of the response styles of five LLMs tasked with annotating quality scores for human reference data from the SummEval dataset Fabbri et al. (2021). We averaged the scores across four metrics for each data point and plotted both the Pearson correlation coefficient ($\rho$) and the average score distributions of the five models. After perturbing the original data by replacing some named entities with fictional ones in the summaries (Fictional Named Entities in Table 1), we repeated the quality evaluation. As shown in Figure 2, all models detected the changes and adjusted their scores accordingly, though their scoring distributions varied significantly. For instance, Llama3 Meta , Mistral Jiang et al. (2023), and Qwen Bai et al. (2023) models assign higher scores to the original data and moderate scores to the perturbed data. In contrast, GPT4-Turbo OpenAI (2023) and Vicuna Chiang et al. (2023) models tend to give moderate scores to the original data and much lower scores to the perturbed data. The variance in the response distributions indicates the presence of bias that can significantly affect alignment ($\rho$), illustrating that alignment is not a direct or credible metric for assessing the ability of LLMs as NLG evaluators. It is crucial to develop a new metric and measurement for evaluation that is not influenced by the evaluators’ biased response styles, ensuring a more accurate and fair assessment of LLM capabilities.
@@ 4 DHP Benchmarking Framework
The overall framework is shown in Figure 3. First, for a specific NLG task, we employ a hierarchical perturbation pipeline to transform high-quality reference data into various forms of lower-quality data. Subsequently, an LLM evaluates both the original and perturbed texts respectively using predefined metrics, generating several sets of rating scores. We then conduct a statistical analysis of these scores. For each pair of scores, original and perturbed, we apply the Wilcoxon Signed-Rank Test to determine the differences in their distributions, achieving this with a confidence level expressed as a $p$-value. This test specifically assesses differences in pairwise scores without focusing on absolute values, thereby minimizing the impact of models’ response styles. Following this, we combine the $p$-values from different metrics, incorporating Expert Weights ($EW$) to tailor the aggregated $p$-values to the specific metrics of the corresponding perturbation methods. These combined $p$-values are then transformed into discernment scores, which serve as a direct measure for assessing and comparing the NLG evaluation capabilities of LLMs for this particular task.
@@ 4.1 Step 1: Hierarchical Perturbation
First, at the character level, we alter some characters or letters in the given $N$ original texts independently. At the word and sentence levels, we degrade the text by processing entire words or sentences, respectively. For NLG tasks involving very short texts, sentence-level perturbation is considered optional. For each level of perturbation, we choose either a rule-based or an LLM-based method, enhancing the diversity of the perturbation’s content and format. Additionally, if the text data is sufficiently long for more perturbation, we implement two degrees of perturbation – minor and major – for each method. These different degrees of perturbation will influence the difficulty that LLMs face in detecting issues within the text. The detailed perturbation methods for each task are shown in Table 1.
@@ 4.2 Step 2: LLM evaluation
After conducting the LLM evaluation on a dataset consisting of $N$ datapoints, we obtain several sets of absolute evaluation scores shown in Figure 3:
@@
where each $\{S\}$ is a set of $N$ evaluation scores. The superscripts $0,1,\ldots,P$ on $S$ represent the original data (00) and the $P$ types of perturbed data ($1,\dots,P$), respectively. The subscripts $m_{1},\ldots,m_{M}$ represent the $M$ different metrics used in the dataset. For instance, in the SummEval dataset Fabbri et al. (2021), there are four evaluation metrics: coherence, consistency, fluency, and relevance.
@@ 4.3.1 Wilcoxon Signed-Rank Test
In our analysis, we adopt a one-sided alternative hypothesis. The resulting $p$-value indicates the confidence level at which we can reject the null hypothesis – that $\{S^{0}_{m_{j}}\}$ and $\{S^{i}_{m_{j}}\}$ have the same distribution – and accept the alternative hypothesis – that $\{S^{0}_{m_{j}}\}$ has a greater distribution than $\{S^{i}_{m_{j}}\}$. We consider a difference to be statistically significant if $p^{i}_{m_{j}}<0.05$. A lower $p$-value represents a more significant score difference between the original data and perturbed data. Totally we can get $P$ sets of $p$-values for the $M$ metrics as shown in Figure 3:
@@ 4.3.2 Harmonic Mean p-value and Expert Weights
To address this, we introduce HMP with Expert Weights ($EW$). We conduct a survey involving $10$ NLP experts who are presented with the specific NLG evaluation tasks and metric definitions. They are asked to identify which metric should be most impacted by different quality problems corresponding to the perturbation methods. These preferences are then aggregated to construct $EW$. For instance, a particular quality issue get votes for “coherence”, “consistency”, and “fluency” are $4,1,$ and $5$, respectively, the $EW$ for the corresponding perturbation would be $[0.4,0.1,0.5]$. The $EW$ makes the combination more targeting on those $p$-values that are highly influenced by the perturbation. This weighting makes the $p$-value combination more targeted, focusing on those metrics most influenced by the perturbation. Consequently, the weighted combined $p$-values offer a more precise measure of the LLM evaluators’ ability to not only detect issues but also correctly assign lower scores to the impacted metrics.
@@ 5 Benchmarking LLM Discernment
We evaluate five series of LLMs with varying parameter sizes: the GPT-series Wang et al. (2023a), which includes GPT3.5-Turbo and GPT4-Turbo; the Llama3-series Meta ; the Vicuna1.5 series Chiang et al. (2023); Mistral-7B Jiang et al. (2023); and the Qwen-series Bai et al. (2023).
@@ 6 Conclusion
We introduce the DHP benchmark to assess the discernment capabilities of LLMs as evaluators across various NLG tasks. Our approach not only provides benchmarking results for LLMs but also establishes a robust framework to evaluate how effectively LLMs can identify quality issues, thus serving as competent NLG evaluators. While most models generally perform well, their performance is significantly influenced by factors such as model size, task type, and dataset complexity. By pinpointing particular weaknesses of LLMs in evaluating NLG tasks, this benchmark aids researchers in improving LLM performance going forward.
@@ 7 Limitations
The limitation of our DHP benchmark is that it generates discernment scores specific to each NLG dataset. A comprehensive assessment of the general evaluation capabilities of LLMs across all NLG tasks remains an open challenge. Additionally, the benchmark’s current focus on English-related text limits its generalizability across different languages and cultural contexts, potentially affecting its reliability in the general multilingual NLG tasks.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly serving as evaluators in Natural Language Generation (NLG) tasks; this is often referred to as “LLM-as-a-judge” paradigm. However, the capabilities of LLMs in evaluating NLG quality remain underexplored. Current studies depend on human assessments and simple metrics that fail to capture the discernment of LLMs across diverse NLG tasks. To address this gap, we propose the Discernment of Hierarchical Perturbation (DHP) benchmarking framework, which provides quantitative discernment scores for LLMs. This framework leverages hierarchically perturbed text data and statistical tests to systematically measure the NLG evaluation capabilities of LLMs. We re-established six evaluation datasets for this benchmark, covering four NLG tasks: Summarization, Story Completion, Question Answering, and Translation. Our comprehensive benchmarking of five major LLM families provides critical insight into their strengths and limitations as NLG evaluators. Our dataset is available at https://huggingface.co/datasets/YCWANGVINCE/DHP_Benchmark.
@@
Yicheng Wang1$\dagger$, Jiayi Yuan2$\dagger$, Yu-Neng Chuang2, Zhuoer Wang1, Yingchi Liu3, Mark Cusick3, Param Kulkarni3, Zhengping Ji 3, Yasser Ibrahim3, Xia Hu2, 1Texas A&M University, 2Rice University, 3Axon Enterprise, Inc.
22footnotetext: Equal Contribution.
@@ 1 Introduction
Large Language Models (LLMs) play a crucial role in the field of Natural Language Generation (NLG), advanced wide real-world applications including education Latif et al. (2023), healthcare Yuan et al. (2023), business Teubner et al. (2023), etc. The strong capabilities of LLMs allow them not only to serve as text generators but also increasingly as powerful evaluators of text quality Chiang and Lee (2023); Liu et al. (2023a); Li et al. (2024). Their role as evaluators is crucial for advancements in various applications, such as summarization, story completion, question answering, and translation Li et al. (2024); Wang et al. (2023); Chuang et al. (2024). LLMs are expected to serve as NLG evaluators, providing reasonable quality scores based on different quality metrics with specially designed evaluation prompts.
@@
(1) Lack of Clear and Unbiased Measurement: There is no clear measurement for the capability of LLM evaluators. Existing methods rely on aligning with human scores Chiang and Lee (2023); Liu et al. (2023a), but these scores themselves are subject to biased response styles Schoch et al. (2020).
(2) Multiple Evaluation Metrics: Evaluating NLG quality requires considering multiple metrics. For example, in summarization tasks, metrics such as coherence, consistency, and fluency are essential considerations Liu et al. (2023a); Fabbri et al. (2021); Gabriel et al. (2021). However, LLMs might struggle with correlations between these metrics Hu et al. (2024), potentially leading to misinterpretation and incorrect scoring, which makes it difficult to assess their effectiveness as evaluators.
To address these challenges, we introduce a novel DHP benchmarking frameworkDiscernment of Hierarchical Perturbation – for quantitatively measuring the evaluation capabilities of LLMs. We propose the concept of discernment scores, systematically derived from hierarchically perturbed text data and statistical tests. Reference data is perturbed using multiple hierarchical methods, and differences in LLM evaluation scores are analyzed using the Wilcoxon Signed-Rank Test Wilcoxon (1945). To obtain more reliable overall evaluation results, harmonic mean $p$-values and expert-assigned weights are applied to integrate multiple metrics. The final $p$-value is then converted into a Discernment Score, providing a quantitative measure of the NLG evaluation capabilities of LLMs. This approach enables a more rigorous and comprehensive assessment of LLM performance, independent of the specific response styles exhibited by the models.
@@
This study re-establishes six evaluation datasets across four key NLG tasks: Summarization, Story Completion, Question Answering, and Translation. Each dataset undergoes hierarchical perturbation and is utilized to challenge the evaluative capabilities of LLMs in distinct ways, providing a robust foundation for benchmarking. The datasets include a range of text perturbation methods, from minor character-level problems to significant sentence-level alterations, enabling a thorough examination of the potential discernment limits of LLMs.
@@
Our comprehensive benchmarking, based on newly defined quantitative discernment scores, is conducted across five major LLM series. This methodology uncovers critical insights into their effectiveness as NLG evaluators and provides a detailed understanding of their performance. This benchmark reveals important trends and patterns in the LLM evaluator capacities, highlighting areas of strength as well as potential shortcomings.
@@
The DHP benchmark aims to fill existing gaps by offering a quantitative framework for assessing LLMs’ evaluation capabilities and emphasizing the necessity of considering multiple metrics for accurate and reliable evaluations. We summarize our contributions as follows.
@@
  • Develop the DHP benchmarking framework, introducing quantitative discernment scores for LLMs as NLG evaluators based on hierarchical perturbation.
@@
  • Re-establish six evaluation datasets across four NLG tasks to evaluate the discernment of LLM evaluators.
@@
  • Benchmark five LLM series to analyze their performance and effectiveness in NLG evaluation.
@@
2 Challenge: Biased Response Styles
@@
We illustrate this with an example of the response styles of five LLMs tasked with annotating quality scores for human reference data from the SummEval dataset Fabbri et al. (2021). We averaged the scores across four metrics for each data point and plotted both the Pearson correlation coefficient ($\rho$) and the average score distributions of the five models. After perturbing the original data by replacing some named entities with fictional ones in the summaries (Fictional Named Entities in Table 1), we repeated the quality evaluation. As shown in Figure 2, all models detected the changes and adjusted their scores accordingly, though their scoring distributions varied significantly. For instance, Llama3 Dubey et al. (2024), Mistral Jiang et al. (2023), and Qwen Bai et al. (2023) models assign higher scores to the original data and moderate scores to the perturbed data. In contrast, GPT4-Turbo OpenAI (2023) and Vicuna Chiang et al. (2023) models tend to give moderate scores to the original data and much lower scores to the perturbed data. The variance in the response distributions indicates the presence of bias that can significantly affect alignment ($\rho$), illustrating that alignment is not a direct or credible metric for assessing the ability of LLMs as NLG evaluators. It is crucial to develop a new metric and measurement for evaluation that is not influenced by the evaluators’ biased response styles, ensuring a more accurate and fair assessment of LLM capabilities.
@@ 3 DHP Benchmarking Framework
The overall framework is shown in Figure 3. First, for a specific NLG task, we employ a hierarchical perturbation pipeline to transform high-quality reference data into various forms of lower-quality data. Subsequently, an LLM evaluates both the original and perturbed texts, respectively, using predefined metrics, generating several sets of rating scores. We then conduct a statistical analysis of these scores. For each pair of scores, original and perturbed, we apply the Wilcoxon Signed-Rank Test to determine the differences in their distributions, achieving this with a confidence level expressed as a $p$-value. This test specifically assesses differences in pairwise scores without focusing on absolute values, thereby minimizing the impact of models’ response styles. Following this, we combine the $p$-values from different metrics, incorporating Expert Weights ($EW$) to tailor the aggregated $p$-values to the specific metrics of the corresponding perturbation methods. These combined $p$-values are then transformed into discernment scores, which serve as a direct measure for assessing and comparing the NLG evaluation capabilities of LLMs for this particular task.
@@ 3.1 Step 1: Hierarchical Perturbation
First, at the character level, we alter some characters or letters in the given $N$ original texts independently. At the word and sentence levels, we degrade the text by processing entire words or sentences, respectively. For NLG tasks involving very short texts, sentence-level perturbation is considered optional. For each level of perturbation, we choose either a rule-based or an LLM-based method, enhancing the diversity of the perturbation’s content and format. Additionally, if the text data is sufficiently long for more perturbation, we implement two degrees of perturbation–minor and major–for each method. These different degrees of perturbation are the difficulty that LLMs face in detecting issues within the text. The detailed perturbation methods for each task are shown in Table 1.
@@ 3.2 Step 2: LLM evaluation
After conducting the LLM evaluation on $N$ datapoints, we obtain several sets of absolute evaluation scores shown in Figure 3:
@@
where each $\{S\}$ is a set of $N$ evaluation scores. The superscripts $0,1,\ldots,P$ on $S$ represent the original data (00) and the $P$ types of perturbed data ($1,\dots,P$), respectively. The subscripts $m_{1},\ldots,m_{M}$ represent the $M$ different metrics used in the dataset. For instance, in the SummEval dataset Fabbri et al. (2021), there are four evaluation metrics, namely: coherence, consistency, fluency, and relevance.
@@ 3.3.1 Wilcoxon Signed-Rank Test
In our analysis, we adopt a one-sided alternative hypothesis. The resulting $p$-value indicates the confidence level at which we can reject the null hypothesis – that $\{S^{0}_{m_{j}}\}$ and $\{S^{i}_{m_{j}}\}$ have the same distribution – and accept the alternative hypothesis – that $\{S^{0}_{m_{j}}\}$ has a greater distribution than $\{S^{i}_{m_{j}}\}$. We consider a difference to be statistically significant if $p^{i}_{m_{j}}<0.05$. A lower $p$-value represents a more significant score difference between the original data and perturbed data. In total, we can get $P$ sets of $p$-values for the $M$ metrics, as shown in Figure 3:
@@ 3.3.2 Harmonic Mean p-value and Expert Weights
To address this, we introduce HMP with Expert Weights ($EW$). We conduct a survey involving $10$ NLP experts who are presented with the specific NLG evaluation tasks and metric definitions. They are asked to identify which metric should be most impacted by different quality problems corresponding to the perturbation methods. These preferences are then aggregated to construct $EW$. For instance, a particular quality issue get votes for “coherence”, “consistency”, and “fluency” are $4,1,$ and $5$, respectively, the $EW$ for the corresponding perturbation would be $[0.4,0.1,0.5]$. The $EW$ makes the combination more targeting those $p$-values that are highly influenced by the perturbation. This weighting makes the $p$-value combination more targeted, focusing on those metrics most influenced by the perturbation. Consequently, the weighted combined $p$-values offer a more precise measure of the LLM evaluators’ ability to not only detect issues but also correctly assign lower scores to the impacted metrics.
@@ 4 Benchmarking LLM Discernment
We evaluate five series of LLMs with varying parameter sizes: the GPT-series Wang et al. (2023); OpenAI (2023), which includes GPT3.5-Turbo and GPT4-Turbo; the Llama3-series Dubey et al. (2024); the Vicuna1.5 series Chiang et al. (2023); Mistral-7B Jiang et al. (2023); and the Qwen-series Bai et al. (2023).
@@ 4.2 Other Observations
5 Related Work
Recent advancements highlight the significant potential of utilizing LLMs as evaluators for a variety of NLP tasks. Extensive empirical evidence supports this viewpoint, as demonstrated by studies Liu et al. (2023a); Chiang and Lee (2023); Hu et al. (2024); Desmond et al. (2024); Wang et al. (2023), which assert that the evaluation behaviors of pretrained LLM-based evaluators are well-aligned with those of human preference Liu et al. (2023b). Liusie et al.Liusie et al. (2024) further show that comparative assessments using LLM evaluators outperform prompt-based techniques, though they identify potential positional biases and propose corresponding solutions. Despite the great assessment performance of a single LLM, advanced studies involve multi-LLM agents Chan et al. (2023); Zhang et al. (2023); Li et al. (2023) or human experts Gao et al. (2024); Li et al. (2024) to further increase the judging capability.
While the application of LLMs as judges is a burgeoning area of research, it is imperative to assess their reliability and effectiveness in evaluative roles. To this end, several benchmarks have been recently proposed to evaluate LLMs as judges. For example, JudgeBench  Tan et al. (2024) is designed to assess LLM-based judges on challenging response pairs spanning knowledge, reasoning, math, and coding. Additionally, LLM-judge-eval  Wei et al. (2024) evaluates tasks such as summarization and alignment, incorporating metrics like flipping noise and length bias.
However, despite the progress in LLMs as judges, several challenges persist. First, human involvement remains a crucial factor in both evaluation and alignment, which raises concerns about the extent to which human biases influence LLM-based evaluations. Second, human evaluators themselves are inherently biased, meaning that even if an LLM aligns well with human preferences, it does not necessarily guarantee fairness or accuracy. Additionally, LLMs may misinterpret NLG evaluation metrics Hu et al. (2024), making simple alignment scores unreliable. To overcome these challenges, our work focuses on developing automated and comprehensive methodologies to test the reliability of LLM-based evaluations.
@@ 6 Conclusion
We introduce the DHP benchmark to assess the discernment capabilities of LLMs as evaluators across various NLG tasks. Our approach not only provides benchmarking results for LLMs but also establishes a robust framework to evaluate how effectively LLMs can identify quality issues, thus serving as competent NLG evaluators. While most models generally perform well, their performance is significantly influenced by factors such as model size, task type, and dataset complexity. By identifying specific weaknesses of LLMs in evaluating NLG tasks, this benchmark aids researchers in enhancing “LLM-as-a-judge” methodologies and improving overall LLM performance.
@@ 7 Limitations
While our DHP benchmark provides a systematic and scalable way to assess LLMs’ ability to detect targeted quality issues, it does not offer a complete picture of how these models perform on every aspect of NLG evaluation. First, the discernment scores are generated on a dataset-by-dataset basis, so a truly comprehensive assessment of LLMs across different NLG tasks remains an open challenge. Next, although our framework is designed to reduce reliance on human annotations, it does not fully replace the depth and contextual insight that human evaluations provide. Our perturbation-driven approach highlights particular types of errors rather than capturing the broad spectrum of real-world NLG complexities. Consequently, DHP is best viewed as a complementary tool, and further work is needed to expand its scope to more diverse tasks, languages, and cultural settings, as well as to integrate human judgment for a more holistic evaluation of LLMs.